OpenStudy (anonymous):

Solve (x/3) + 2 > 8.

OpenStudy (anonymous):

\[\frac{x}{3}+2>8 \Rightarrow \frac{x}{3}+2-2>8-2 \Rightarrow 3\cdot \frac{x}{3}>3\cdot 6\] And then we have \[x>18\]